OpenStudy (anonymous):

Evaluate the following limits algebraically: lim x -> 0 1/3+x - 1/3/x lim x -> infintity 7x^3 + 5x/(8x^9 + x^6 + 4)^ 1/3

OpenStudy (mathmath333):

is it? \(\huge\tt \color{black}{\lim_{x\to0}\frac{1}{3+x}-\frac{1}{\frac{3}{x}}}\)

OpenStudy (anonymous):

no i should have used parenthesis lim x -> (1/3+x - 1/3) / x

OpenStudy (mathmath333):

\(\large\tt \color{black}{\lim_{x\to0}\dfrac{\dfrac{1}{3+x}-\dfrac{1}{3}}{x}}\) \(\large\tt \color{black}{\lim_{x\to0}\dfrac{1}{x}\times(\dfrac{1}{3+x}-\dfrac{1}{3})}\) \(\large\tt \color{black}{\lim_{x\to0}\dfrac{1}{x}\times(\dfrac{3}{3(3+x)}-\dfrac{(x+3)}{3(x+3)})}\) \(\large\tt \color{black}{\lim_{x\to0}\dfrac{1}{x}\times(\dfrac{3-x-3}{3(3+x)})}\) \(\large\tt \color{black}{\lim_{x\to0}\dfrac{1}{x}\times\dfrac{-x}{3(3+x)}}\) \(\large\tt \color{black}{\lim_{x\to0}\dfrac{-1}{3(3+x)}}\) \(\large\tt \color{black}{=\dfrac{-1}{3(3+0)}}\) \(\large\tt \color{black}{=\dfrac{-1}{9}}\)

OpenStudy (mathmath333):

this one is correct

OpenStudy (anonymous):

thank you so much!!

OpenStudy (mathmath333):

yw